OpenStudy (anonymous):

If m = 35 cm and n = 37 cm, what is the length of l?

OpenStudy (anonymous):

What is "l" here?

OpenStudy (anonymous):

You have studied about Pythagorean Theorem??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So, try to use that..

OpenStudy (anonymous):

problem is that i don't remember anything about it..

OpenStudy (anonymous):

Oh.. Pythagorean Theorem says that: In right angled triangled, if you have two sides and a hypotenuse, then : \[(hypotenuse)^2 = (first \; \; side)^2 + (second \; \; side)^2\]

OpenStudy (anonymous):

so is this case, what would "l" be?

OpenStudy (anonymous):

"l" is your first or second side, you can take what you want.. Hypotenuse = n First side = m Second Side = l

OpenStudy (anonymous):

"l" is your one of the sides..

OpenStudy (anonymous):

Try to use that..

OpenStudy (anonymous):

Using Pythagoras Theorem: \[n^2 = l^2 + m^2\]

OpenStudy (anonymous):

You are given with m and n values.. Try to find "l" here..